OpenStudy (anonymous):

Which of the following is a polynomial with roots –2, –3i, and 3i ? x3 + 4x2 + 9x + 24 x3 – 4x2 + 9x – 24 x3 + 2x2 + 9x + 18 x3 – 2x2 + 9x – 18

OpenStudy (amoodarya):

f(x)=(x-(-2)) *(x-3i) *(x+3i) = (x+2)(x^2+9) so you can go on

OpenStudy (anonymous):

x3+2x2+9x+18

OpenStudy (amoodarya):

x3 + 2x2 + 9x + 18 = x^2 (x+2) +9(x+2) is another way